OpenStudy (anonymous):

Explain, in complete sentences, when you would use an open circle or a closed circle, and when you would shade to the right or left, to graph an inequality on the number line

OpenStudy (lgbasallote):

you shade the circle if the point is included (when there is \(\ge\) or \(\le\)) you use open circle when it's neither of those (i.e. \(\ne\), < or >) you shade to the right when x is greater than (i.e. \(\ge\) or >) you shade to the left when x is less than (i.e. \(\le\) or <)

OpenStudy (lgbasallote):

do you get that?